OpenStudy (loser66):

Write the negation of this statement, avoiding use of the word "not" "For all real numbers x, there exists a positive integer N such that \(N-1\leq x^2\leq N\) " Please, help

OpenStudy (phi):

we have to use https://en.wikipedia.org/wiki/There_exists#Negation

OpenStudy (loser66):

Thanks a ton. :)

OpenStudy (loser66):

Ok, check my work please I symbolize the statement as \[\forall x \in R, \exists~~ N >0~~ | N-1\leq x^2 \leq N\] \[\neg(\forall x \in R, \exists~~ N >0~~ | N-1\leq x^2 \leq N)\\\exists x \in R, \forall N> 0 ~~ | N-1 \geq x^2 \cup x^2\geq N\] Verbalize it: "There exists a real number x, for all positive N such that x^2 is lesser or equal N-1, or x^2 greater or equal N" Am I right?? Please check.

OpenStudy (anonymous):

there exist a real number x , such that for all possitive N N-1>x^2 or x^2 >N so your correct ^^

OpenStudy (loser66):

Thank you @ikram I made mistake at \(\geq\) . It should be > only. hihiihi...